# Absence of antisymmetric tensor fields: Clue from Starobinsky model of $f(R)$ gravity

### v3

@@ 5.1 Case I: $n=2$, Starobinsky model
$\frac{d\lambda}{dN}=\sqrt{\frac{3}{2}}\lambda^{2}x\left[1-\frac{1}{\lambda}\sqrt{\frac{8}{3}}\right]$ (5.1)
Note that the other equations in (3.15) remain the same. We solve the above equations numerically, with suitable initial conditions such that radiation-matter equality happens around redshift 3300 and matter in the present universe constitutes approximately 30% of the energy density. In Fig (2), we present different density parameters, scalar field $\tilde{\phi}$ and the equation of state $\omega_{\tilde{\phi}}=\frac{\tilde{P}_{\tilde{\phi}}}{\tilde{\rho}_{\tilde{\phi}}}$ with respect to $N(=\ln{\tilde{a}})$ The density parameters of different components mimic the behaviours in standard model of cosmology. From the plot it is evident that the energy density of $\tilde{\phi}$ field never dominates the evolution until recently.
Figure 2: Evolution of density parameters, scalar field $\tilde{\phi}$ and $\omega_{\tilde{\phi}}$ in Starobinsky model
$\tilde{\phi}$ remains positive along the entire evolution and does not vary much. $\omega_{\tilde{\phi}}$ remains constant for a long period and starts varying late, around $z\sim 7$ and finally approaches $-1$ at present. So, in the late time the field behaves as dark energy. This behaviour of $\omega_{\tilde{\phi}}$ resembles the equation of state of the thawing quintessence field.
We further investigate the potential and find that the scalar field rolls down the potential $V(\tilde{\phi})$ but does not reach the minima. This is due to the coupling between the scalar field and matter (see equations (3.11) and (3.12)). Plotting both the terms separately, we realised that the potential arising due to coupling with matter $V_{m}$ is far more dominant than the potential arising from modified gravity $V(\tilde{\phi})$. Hence the scalar field did not reach the minima. In Fig (3), we present the three different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$.
@@ 5.2 Case 2: $n\neq 2$, $f(R)=R+\alpha_{n}R^{n}$
$\frac{d\lambda}{dN}=\sqrt{\frac{3}{8}}x(4\sqrt{8/3}-6\lambda)(\sqrt{8/3}-\lambda)$ (5.2)
along with the other equations in equation(3.15). Following similar methods as mentioned above, we solve the equations numerically, with suitable initial conditions. We did a similar analysis of the cosmological parameters as we have done in the case of $n=2$. Though the form of the potential of $V(\phi)$ in this case is different from the Starobinsky model,interestingly, we find that the behavior of the density parameters and the equation of state almost exactly matches the Starobinsky model over the entire period of evolution (Fig 4). This is because here also the potential $V_{m}$ arising due to matter coupling dominates, hence the effective potential $V_{eff}$ almost similarly matches the case of $n=2$ (Fig 5). The only difference is that the scalar field is negative for the entire period as it rolls down the potential. For all other values of n, ($\neq 2$), we get the same results.
@@ 5.3 $f(R)=R+\alpha_{n}R^{n}$ models with Cosmological Constant
We have also carried out the entire study including the Cosmological Constant $\Lambda$ along with the other matter components in the observable universe i.e, $\tilde{\rho}_{M}=\tilde{\rho}_{r}+\tilde{\rho}_{m}+\tilde{\rho}_{\Lambda}$. The scalar field $\tilde{\phi}$ also couples with $\Lambda$ also, apart from matter. So, equations (3.10) and (3.11) changes to include $\Lambda$, while (3.12) and (3.13) remain unchanged. Also, we have one more coupled equation for $\Lambda$.
$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}+\tilde{\rho}_{\Lambda}\bigg{]}$ (5.3)
$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m}+4\tilde{\rho_{\Lambda}})\tilde{\phi}^{\prime}$ (5.4)
$\tilde{\rho}_{\Lambda}^{\prime}=4\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{\Lambda}\tilde{\phi}^{\prime}$ (5.5)
@@
$\frac{dx}{dN}$ $=\frac{x}{2}(3x^{2}+3-3y^{2}+r-3l)-3x+\sqrt{\frac{3}{2}}\lambda y^{2}$ (5.6)
@@
We perform a similar numerical analysis with the new set of equations for the potential given by (4.1) with different $n$ as has been done without $\Lambda$. Surprisingly, we found that inclusion of the Cosmological constant does not effect the evolution of different components. And like the earlier scenario, the coupling term with matter $V_{m}$ is far more dominant than the potential $V(\tilde{\phi})$ arising out of modified gravity as well as the coupling with the cosmological constant $V_{l}$. Hence driven mostly by the matter coupling term $V_{m}$, the evolution of different components remain similar. The value of the scalar field for different $n$ also matches our earlier findings. For $n=2$, scalar field remains positive throughout, while for other models $(n\neq 2)$ it remains negative even after inclusion of $\Lambda$. Here we present only one case $n=2$ as an example.
Figure 6: Evolution of different cosmological parameters and potentials including $\Lambda$ for $n=2$
@@ 6 Results
We have considered a very general class of modified gravity of the form $f(R)=R+\alpha_{n}R^{n}$ gravity, which is equivalent to Einstein gravity along with a scalar degree of freedom $\tilde{\phi}$ following the potential given by (4.1), when conformally transformed in Eintein frame. Depending on the behaviour of the potential we divided our cases in two categories $n=2$ (Case I) and $n\neq 2$ (Case 2). We investigated the evolution of different components like scalar field, radiation and matter in a background isotropic and homogeneous universe represented by FRW metric. After the whole analysis we found that in case $n=2$, we found that $\tilde{\phi}$ remains positive along the entire evolution and does not vary much. The matter - scalar field coupling does not allow the scalar field to reach the minima at zero. Such values of $\tilde{\phi}$ for the last two terms in action (2.4) provide an additional exponential suppression. This makes the signature of the antisymmetric fields (both rank 2 and 3) heavily suppressed thereby failing to leave some imprints in the observable universe. The investigations for the case of $n\neq 2$ reveal $\tilde{\phi}$ to be negative all along (again matter-$\tilde{\phi}$ coupling not allowing $\tilde{\phi}$ to reach zero). Such values of $\tilde{\phi}$ would imply enhancement of the coupling for the last two terms in action (2.4) by an exponential factor. This indicates that signatures of the antisymmetric fields of rank 2 and 3 should be present, which is contrary to the observation. For all other values of $n\neq 2$, we get the same results. We also performed the analysis adding Cosmological constant $\Lambda$ along with the earlier components in the FRW background. Even though the equations change significantly, the nature of the scalar field for different $n$ matches our earlier findings. So with or without $\Lambda$, the antisymmetric fields of rank 2 and 3 gets heavily suppressed in case of Starobinsky model $(n=2)$, justifying the invisibility of these fields, while their couplings gets enhanced for other cases $(n\neq 2)$, implying presence of such fields, contrary to the observation.
@@ 7 Discussion and Conclusion
In our current universe, we don’t see any signs of higher rank antisymmetric tensor fields having an impact on natural events. We explain the suppression of these fields in the light of the higher curvature $f(R)$ theory. It has been extensively discussed in literature that modified $f(R)$ gravity when expressed in the Einstein frame, is equivalent to Einstein gravity along with a scalar degree of freedom $(\tilde{\phi})$ following a potential, governed by the form of $f(R)$. Thus, in the Einstein frame, $\tilde{\phi}$ couples with all the other fields through the conformal transformation of the metric. In this work, we first studied how different components like scalar field, radiation and matter evolve in a background isotropic and homogeneous universe represented by FRW metric. Further, we explore how the nature of the scalar field influences the coupling of the massless modes of higher rank antisymmetric fields present in the action (2.4). The coupling, in turn, is responsible for the observation of the possible imprints of massless modes of these antisymmetric fields.
For this study, we have considered the form of $f(R)$ to be $R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has suitable dimensions. As shown in fig(1), the shape of the potential $V(\tilde{\phi})$ for this form is same for all values of $n$, apart from $n=2$. For $n\neq 2$, the potential is zero for positive $\tilde{\phi}$ and increases exponentially for $\tilde{\phi}<0$. The $n=2$ case, Starobinsky model, is different from the rest, as the potential increases and soon reaches a constant for positive $\tilde{\phi}$ but for negative $\tilde{\phi}$, the behaviour is the same as other cases. The field equations with different kinds of matter and scalar field in the homogeneous and isotropic background are coupled nonlinear differential equations. We rewrite the equations as a set of autonomous system of first order differential equations to tackle the complexity. The autonomous system is solved numerically, with suitable initial conditions such that, following standard model of cosmology, the matter radiation equality occurs around redshift $z\sim 3300$ and the matter content is $30\%$ of the total in the present universe.
@@
Most interestingly, for different values of $n$, even though the different matter components, namely matter, radiation and scalar field evolve in the similar fashion with time, $\tilde{\phi}$ is positive only for $n=2$ and negative for $n\neq 2$, throughout. The value of $\tilde{\phi}$ depends on the slope of the potential. For $n=2$, $\tilde{\phi}$ remains positive as it rolls towards the minima ($\tilde{\phi}=0$). Where as, for $n\neq 2$, $\tilde{\phi}$ can roll only over negative value, as on the positive side the potential is zero. However, for both the cases $\tilde{\phi}$ can never reach zero due to coupling with matter. Hence even though both the categories agree with standard model of cosmology, only for Starobinsky model $(n=2)$, the last two terms of the action (2.4) get an extra exponential suppression. For $n\neq 2$, we obtain the enhancement of the coupling term instead of the suppression. As there has been no experimental evidence of the footprint of antisymmetric fields on the present Universe, so $n=2$ i.e., the Starobinsky model is the only model in a general class of modified gravity where $f(R)=R+\alpha_{n}R^{n}$, to support the suppression of such fields. We have also added cosmological constant $\Lambda$ as another component. Even though $\Lambda$ also couples with $\tilde{\phi}$, the effect is very less with respect to matter-$\tilde{\phi}$ coupling which remains the dominant term in the whole analysis. Hence, the dynamics and inference regarding the suppression of the antisymmetric fields remain unchanged. So with or without $\Lambda$ the Starobinsky model is the only model to justify the absence of signature of the massless modes of antisymmetric fields in a general class of modified gravity model.

### v4

@@ 5.1 Case I: $n=2$, Starobinsky model
$\frac{d\lambda}{dN}=\sqrt{\frac{3}{2}}\lambda^{2}x\left[1-\frac{1}{\lambda}\sqrt{\frac{8}{3}}\right]$ (21)
Note that the other equations in (19) remain the same. We solve the above equations numerically, with suitable initial conditions at the epoch of nucleosynthesis at $z\sim 10^{8}$. At such an early epoch radiation is the most dominating component compared to others . We have chosen $x=0.009,~{}y=2.2\times 10^{-15},~{}\lambda=-6\times 10^{-4}~{}{\text{and}}~{}r=0.9999$. Following these conditions radiation-matter equality happens around redshift 3300 and matter in the present universe constitutes approximately 30% of the energy density. In Fig (2), we present different cosmological parameters and the scalar field as they vary with redshift. The left panel presents different density parameters with respect to $\ln(1+z)$, The density parameters of matter and radiation follow their behavior of standard cosmology, while the density parameter of $\tilde{\phi}$ field never dominates the evolution until recently. To further understand the behavior of $\tilde{\phi}$, in the middle panel we plot $\omega_{\tilde{\phi}}(=\frac{\tilde{P}_{\tilde{\phi}}}{\tilde{\rho}_{\tilde{\phi}}})$ with respect to $\ln(1+z)$. $\omega_{\tilde{\phi}}$ remains constant for a long period and starts varying late and finally approaches $-1$ at present. So, in the late time the field behaves as dark energy. This behaviour of $\omega_{\tilde{\phi}}$ resembles the equation of state of the thawing quintessence field. However, since our main goal is to study the nature of the scalar field, in the right most panel we present $\tilde{\phi}$ with respect to $\ln(1+z)$. We observe that $\tilde{\phi}$ remains positive along the entire evolution and does not vary much.
Figure 2: Evolution of density parameters, scalar field $\tilde{\phi}$ and $\omega_{\tilde{\phi}}$ in Starobinsky model(n=2).
To understand the behaviour of the scalar field better, we investigate the potential $V(\tilde{\phi})$ and find that the scalar field rolls down the potential $V(\tilde{\phi})$ but does not reach the minima. This is due to the coupling between the scalar field and matter present on the right hand side of equation (10). Denoting $\frac{\partial V_{m}}{\partial\phi}=\kappa Q\tilde{\rho}_{m}$, one can write the potential $V_{m}$ due to the coupling term. Hence the scalar field rolls down the effective potential $V_{eff}(=V(\tilde{\phi})+V_{m}$). Plotting both the terms $V(\tilde{\phi})$ and $V_{m}$ separately, we realized that $V_{m}$ is far more dominant than $V(\tilde{\phi})$, the potential arising from modified gravity. Hence the scalar field rolls down the effective potential $V_{eff}(\approx V_{m})$ and did not reach the minima of $V(\tilde{\phi})$(see Fig (1)). In Fig (3), we present all the three different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$.
@@ 5.2 Case II: $n\neq 2$, $f(R)=R+\alpha_{n}R^{n}$
$\frac{d\lambda}{dN}=\sqrt{\frac{3}{8}}x(4\sqrt{8/3}-6\lambda)(\sqrt{8/3}-\lambda)$ (22)
along with the other equations in equation(19). Following similar methods as mentioned above, we solve the equations numerically, with suitable initial conditions at the epoch of nucleosynthesis. We set the initial conditions similar to the previous case ($x=0.008,~{}y=1.3\times 10^{-16},~{}\lambda=1.7~{}{\text{and}}~{}r=0.0.99997$ at $\ln(1+z)=18.8$. Note that negative initial values for $\lambda$ are not permissible in this case due to the form of potential.) We did a similar analysis of the cosmological parameters as we have done in the case of $n=2$. Though the form of the potential of $V(\phi)$ in this case is different from the Starobinsky model, interestingly, we find that the behavior of the density parameters and the equation of state are similar to the Starobinsky model over the entire period of evolution (Fig 4). The only difference is that the scalar field is negative for the entire period as it rolls down the potential. Further investigations on the modified gravity potential $V(\phi)$ and coupling potential $V_{m}$ reveal $V_{m}$ dominates over $V(\phi)$ like the previous case. Hence the scalar field rolls down the effective potential $V_{eff}(=V(\tilde{\phi})+V_{m})$ which is $\approx V_{m}$ (Fig 5). So, even though the form of the potential of $V(\phi)$ is different from the Starobinsky model, the cosmological parameters behaves similarly, but the scalar field remains negative. For all other values of n, ($\neq 2$), we get the same results.
@@ 5.3 $f(R)=R+\alpha_{n}R^{n}$ models with Cosmological Constant
We have also carried out the entire study including the Cosmological Constant $\Lambda$ along with the other matter components in the observable universe i.e, $\tilde{\rho}_{M}=\tilde{\rho}_{r}+\tilde{\rho}_{m}+\tilde{\rho}_{\Lambda}$. The scalar field $\tilde{\phi}$ also couples with $\Lambda$ also, apart from matter. So, equations (14) and (15) changes to include $\Lambda$, while (16) and (17) remain unchanged. Also, we have one more coupled equation for $\Lambda$.
$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}+\tilde{\rho}_{\Lambda}\bigg{]}$ (23)
$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m}+4\tilde{\rho_{\Lambda}})\tilde{\phi}^{\prime}$ (24)
$\tilde{\rho}_{\Lambda}^{\prime}=4\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{\Lambda}\tilde{\phi}^{\prime}$ (25)
@@
$\frac{dx}{dN}$ $=\frac{x}{2}(3x^{2}+3-3y^{2}+r-3l)-3x+\sqrt{\frac{3}{2}}\lambda y^{2}$ (26)
@@
We perform a similar numerical analysis with the new set of equations for the potential given by (20) with different $n$ as has been done without $\Lambda$. Surprisingly, we found that inclusion of the Cosmological constant does not effect the evolution of different components. Due to the inclusion of $\Lambda$, the scalar field couples with both matter and $\Lambda$ (equation (24)). So we have two coupling terms $V_{m}$ and $V_{l}$. But, even then, like the earlier scenario, $V_{m}$ is more dominant than both $V_{l}$ and $V(\tilde{\phi})$ of modified gravity. Hence driven mostly by the matter coupling term $V_{m}$, the evolution of different components remain similar. The value of the scalar field for different $n$ also matches our earlier findings. For $n=2$, scalar field remains positive throughout, while for other models $(n\neq 2)$ it remains negative even after inclusion of $\Lambda$. Here we present only one case $n=2$ as an example.
Figure 6: Evolution of different cosmological parameters and potentials including $\Lambda$ for $n=2$. The initial conditions in these plots are same as mentioned in section 5.1 along with $l=6\times 10^{-29}$ at $\log(1+z)=18.8$ for the new variable $l$.
@@ 6 Results
We have considered a very general class of modified gravity of the form $f(R)=R+\alpha_{n}R^{n}$ in an early universe around the epoch of nucleosynthesis. When conformally transformed in Eintein frame, modified gravity is equivalent to Einstein gravity along with a scalar degree of freedom $\tilde{\phi}$ following the potential given by (20). Depending on the behaviour of the potential we divided our cases in two categories $n=2$ (Case I) and $n\neq 2$ (Case II). We investigate the evolution of different components like scalar field, radiation and matter in a background isotropic and homogeneous universe represented by FRW metric. After the whole analysis we found that in the case of $n=2$, $\tilde{\phi}$ remains positive along the entire evolution and does not vary much. The matter - scalar field coupling does not allow the scalar field to reach the minima at zero. Such values of $\tilde{\phi}$ for the last two terms in action (4) provide an additional exponential suppression. This makes the signature of the antisymmetric fields (both rank 2 and 3) heavily suppressed thereby failing to leave some imprints in the observable universe. The investigations for the case of $n\neq 2$ reveal $\tilde{\phi}$ to be negative all along (again matter-$\tilde{\phi}$ coupling not allowing $\tilde{\phi}$ to reach zero). Such values of $\tilde{\phi}$ would imply enhancement of the coupling for the last two terms in action (4) by an exponential factor. This indicates that signatures of the antisymmetric fields of rank 2 and 3 should be present, which is contrary to the observation. For all other values of $n\neq 2$, we get the same results. We also performed the analysis adding Cosmological constant $\Lambda$ along with the earlier components in the FRW background. Even though the equations change significantly, the nature of the scalar field for different $n$ matches our earlier findings. So with or without $\Lambda$, the antisymmetric fields of rank 2 and 3 gets heavily suppressed in case of Starobinsky model $(n=2)$, justifying the invisibility of these fields, while their couplings gets enhanced for other cases $(n\neq 2)$, implying presence of such fields, contrary to the observation.
@@ 7 Discussion and Conclusion
In our current universe, we don’t see any signs of higher rank antisymmetric tensor fields having an impact on natural events. We explain the suppression of these fields in the light of the higher curvature $f(R)$ theory. It has been extensively discussed in literature that modified $f(R)$ gravity when expressed in the Einstein frame, is equivalent to Einstein gravity along with a scalar degree of freedom $(\tilde{\phi})$ following a potential, governed by the form of $f(R)$. Thus, in the Einstein frame, $\tilde{\phi}$ couples with all the other fields through the conformal transformation of the metric. In this work we study in an early universe post reheating, how different components of energy evolve in a background isotropic and homogeneous universe represented by FRW metric. Further, we explore how the nature of the scalar field influences the coupling of the massless modes of higher rank antisymmetric fields present in the action (4). The coupling, in turn, is responsible for the observation of the possible imprints of massless modes of these antisymmetric fields.
For this study, we have considered the form of $f(R)$ to be $R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has suitable dimensions. As shown in fig(1), the shape of the potential $V(\tilde{\phi})$ remain same for all values of $n\neq 2$ the potential is zero for positive $\tilde{\phi}$ and increases exponentially for $\tilde{\phi}<0$. But, the $n=2$ Starobinsky model is different from others, as the potential is positive for positive $\tilde{\phi}$ and negative for negative $\tilde{\phi}$ with a minima at $\tilde{\phi}=0$. The field equations with different kinds of matter and scalar field in the homogeneous and isotropic background are coupled nonlinear differential equations. We rewrite the equations as a set of autonomous system of first order differential equations to tackle the complexity. The autonomous system is solved numerically, with suitable initial conditions at early epoch of nucleosynthesis.
@@
Most interestingly, for different values of $n$, even though the behaviour of the density parameters of different components, namely matter, radiation and scalar field remain similar, $\tilde{\phi}$ is positive only for $n=2$ and negative for $n\neq 2$, throughout. The value of $\tilde{\phi}$ depends on the slope of the potential. For $n=2$, $\tilde{\phi}$ remains positive as it rolls towards the minima ($\tilde{\phi}=0$) on the positive side of the potential. Where as, for $n\neq 2$, $\tilde{\phi}$ can roll only over negative value, as on the positive side the potential is zero. However, for both the cases $\tilde{\phi}$ can never reach zero due to coupling with matter. Hence even though both the categories agree with standard model of cosmology, only for Starobinsky model $(n=2)$, the last two terms of the action (4) get an extra exponential suppression. For $n\neq 2$, we obtain the enhancement of the coupling term instead of the suppression. As there has been no experimental evidence of the footprint of antisymmetric fields on the present Universe, so $n=2$ i.e., the Starobinsky model is the only model in a general class of modified gravity where $f(R)=R+\alpha_{n}R^{n}$, to support the suppression of such fields. We have also added cosmological constant $\Lambda$ as another component. Even though $\Lambda$ also couples with $\tilde{\phi}$, the effect is very less with respect to matter-$\tilde{\phi}$ coupling which remains the dominant term in the whole analysis. Hence, the dynamics and inference regarding the suppression of the antisymmetric fields remain unchanged. So with or without $\Lambda$ the Starobinsky model is the only model to justify the absence of signature of the massless modes of antisymmetric fields in a general class of modified gravity model.